The cost of manufacturing x tennis balls is C(x)=0.1x^(2)+25x+1000 if 0<x<700 . a) Find the average cost function C(x) . b) Identify all intervals on which C(x) is increasing. c) Identify all intervals on which C(x) is decreasing. d) Identify the x-coordinate for all local extrema of C(x) .
Answer #1

c(x)=0.100²+250 +1000 a) Average cost 0.186+258 Hooo 0.10+25 + 1000 L b) c(oc) is c(x)=o10² +250 +1000 c(x)= 0.20 +25 Hote